The current price of a stock is $ 46.40 and the annual risk-free rate is 4.2 percent. A put option with an exercise price of $55 and one year until expiration has a current value of $ 6.03 . What is the value of a call option written on the stock with the same exercise price and expiration date as the put option? Show your answer to the nearest .01. Do not use $ or , in your answer. Because of the limitations of WEBCT random numbers, some of the options may be trading below their intrinsic value. Note, the given interest rate is an effective rate, so for calculation purposes, you need only discount the using the risk free rate, no e x adjustment is needed.
Answer #1

As per Put Call Parity,

C + (X)e-rt = P + S0

C = Price of Call Option

P = Price of Put Option

S0 = Spot Price of Underlying

X = Exercise price of Option

r = risk free rate

t = time period to expiration

C + (55)e-0.042*1 = 6.03 + 46.40

C = 52.43 - 52.74

C = $-0.31

So,

Price of Call Option = $0.31